# Energy-Efficient Flying LoRa Gateways: A Multi-Agent Reinforcement Learning Approach

### v2

@@ V-B Simulation Results
In Fig. 3(a), we illustrate the association between EDs and GWs in a 3D plane. As shown in the figure, we employ the one-to-many matching scheme from [17], where each GW is associated with multiple LoRa EDs. Consequently, the quota constraint is maintained, and the network load is balanced across GWs. In Fig. 3(b), the training performance is evaluated for four different learning rates ($\alpha$). In this setup, we fixed the clipping range at $\epsilon=0.2$, with 60 EDs and 5 GWs. Each training curve in the figure represents the global cumulative reward over environment timesteps for a specific learning rate: $\alpha=0.0005$, $\alpha=0.0003$, $\alpha=0.003$, and $\alpha=0.03$. It can be observed that $\alpha=0.0003$ achieves the highest rewards with stable convergence throughout the training phase. In contrast, $\alpha=0.03$ performs the worst, yielding very low global cumulative rewards with an unstable training process at different timesteps. Although $\alpha=0.0005$ converges slightly faster than $\alpha=0.0003$, it becomes unstable toward the end of training. Therefore, we utilize $\alpha=0.0003$ as it achieves the highest reward while maintaining stable learning.
In Fig. 3(c), the convergence behavior is examined for different numbers of GWs while keeping EDs fixed at 60, $\epsilon=0.2$, and $\alpha=0.0003$. The results show that with fewer GWs (e.g., 2), the rewards converge faster and stabilize earlier but fail to achieve the highest global system EE. On the other hand, as the number of GWs increases (e.g., 6 and 8), the convergence rate decreases. However, it can be observed that a higher number of GWs leads to greater rewards. This is expected, as more GWs provide better spatial coverage and improved connectivity, resulting in more efficient data aggregation and reduced communication overhead. Consequently, this enhances the overall global system EE and enables better resource allocation.
@@
Figure 3: (a) 3D plane of LoRa ED-gateway association, (b) Training reward with different learning rates $\alpha$, (c) Training reward with different gateways.
In Fig. 4, we compare the optimal EE versus the number of active EDs across multiple multi-agent RL benchmarks, including COMA, MAA2C, and VDN. Our proposed framework consistently achieves the highest global EE due to its CTDE paradigm, which enables efficient EE allocation across the network. Unlike COMA, which suffers from high variance and scalability issues, MAPPO ensures balanced reward distribution. On the other hand, MAA2C outperforms COMA and VDN, its cooperative learning limitations result in declining EE with an increasing number of EDs. Similarly, VDN exhibits the lowest EE due to its lack of explicit agent coordination. Additionally, with five fixed GWs, our approach maintains superior global system EE, whereas other algorithms struggle in denser networks due to coordination failures and communication overhead. It is important to note that as the number of EDs increases, the global EE naturally decreases. A higher number of EDs leads to a higher total TP consumption, increased interference, and reduced data rates, all of which contribute to lower EE. Despite these challenges, MAPPO mitigates this degradation more effectively than the other approaches. In comparison with the closest-performing algorithm (MAA2C), our approach improves the global system EE by 30.5%, 7.2%, 6.5%, 10.9%, and 11.0% for 20, 30, 40, 50, and 60 EDs, respectively.
Figure 4: Total energy efficiency under different number of end-devices
@@ VI Conclusion
In this paper, we considered the joint optimization of assigning spreading factor, transmission power, bandwidth, and ED associations between multiple EDs and GWs while considering resource constraints and the A2G propagation to optimize the global system EE in a multi-flying LoRa network. We showed that the resulting sequential decision-making problem can be modeled as a POMDP, and we proposed a model-free RL algorithm that leverages a novel MAPPO scheme. We used simulation to show that our proposed approach can learn a good policy approximation by optimizing the global system EE in LoRa networks. Interesting directions of future work include integrating trajectory optimization for multiple flying LoRa gateways for efficient resource management. Our approach could be extended by exploring the Age of Information (AoI) scheme to enhance the freshness of data collection.

### v3

@@ V-B Simulation Results
In Fig. 3(a), we present the association between EDs and GWs using a one-to-many matching association scheme, with a utility function designed to maximize the SNR. As illustrated in the figure, the quota constraint for upper-layer nodes (GWs) is respected, and all LoRa EDs are associated with at least one GW in such a way that the overall SNR of the network is maximized.
In Fig. 3(b), we depict the energy efficiency of the studied framework against environment timesteps for varying numbers of LoRa EDs, while keeping the number of GWs fixed at two. The figure shows that the agents experience different network topologies involving dynamic ED associations, mobility, and communication parameters during training. It can be observed that energy efficiency generally increases as the training steps evolve, with most network topologies reaching convergence around $1\times 10^{6}$ timesteps. The agents successfully learn optimal policies for allocating SF, TP, and W, considering both energy efficiency and the A2G wireless link. It is important to highlight that networks with lower ED density achieve higher energy efficiency, whereas those with higher ED density experience reduced energy efficiency. This outcome is logical given the increased complexity and higher overall power consumption introduced by denser network conditions.
In Fig. 3(c), we plot the energy efficiency against training timesteps for various network topologies. For example, we fixed 60 ED positions and varying numbers of GWs. As expected, configurations with a greater number of GWs converge faster and achieve higher energy efficiency, whereas topologies with fewer GWs converge more slowly. This trend is explained by the increased flexibility for EDs to associate with the nearest GW when more GWs are present. Such associations minimize transmission power required during data collection, simultaneously adhering to quota constraints and maximizing the utility function.
@@
Refer to caption Refer to caption Refer to caption
(d) (e) (f)
Figure 3: (a) 3D plane of LoRa ED-GW association, (b) energy efficiency comparison with varying active EDs, (c) EE comparison with varying active GWs, (d) Cumulative power consumption with varying active EDs, (e) Training reward with different MARL benchmarking, (f) energy efficiency comparison with MARL and Non-MARL benchmarking.
Fig. 3(d) further evaluates the power consumption of the studied approach with 2 fixed GWs and a varying number of EDs. As illustrated, network topologies with higher ED densities initially consumed larger amounts of power, slightly above 1 watt. The increased initial consumption is logical due to the greater number of devices requiring communication. However, as training progresses, the agents effectively learn new policies that minimize power usage, thereby significantly improving the overall energy efficiency of the system. In contrast, topologies with fewer EDs consumed only a small amount of power from the start. Specifically, the zoomed-in plot shown in Fig. 3(d) clearly demonstrates the achievement of more realistic and efficient policies across all tested LoRa network topologies.
In Figs. 3(e), we compare the performance of our studied approach against other popular MARL algorithms. First, the cumulative reward functions are plotted against environment timesteps for DRL-based algorithms in Fig.3(e), as this metric applies exclusively to DRL-based approaches. The figure clearly indicates that our proposed method consistently outperforms state-of-the-art MARL algorithms, demonstrating both faster convergence and greater stability. Algorithms such as QMIX fail to converge to optimal policies due to the highly dynamic nature of the environment studied. Additionally, MAA2C exhibits a steadily increasing reward trend but does not converge to an optimal policy within the observed training period.
Furthermore, Fig. 3(f) illustrates a comparison of energy efficiency values against the number of EDs for DRL-based algorithms and other conventional allocation schemes. Our method consistently surpasses both the DRL-based approaches (MAA2C, COMA, and QMIX) and conventional allocation schemes (genetic algorithm and distance-based allocation). The significantly poorer performance of conventional schemes, particularly as the number of EDs increases, is logically consistent with their limited capability in handling complex, high-dimensional, and dynamically changing environments, as exemplified by our proposed LoRa system.
@@ VI Conclusion
In this paper, we considered the joint optimization of assigning spreading factor, transmission power, bandwidth, and ED associations between multiple EDs and GWs while considering resource constraints and the A2G propagation to optimize the global system energy efficiency in a multi-flying LoRa network. We showed that the resulting sequential decision-making problem can be modeled as a POMDP, and we proposed a model-free RL algorithm that leverages a novel MAPPO scheme. We used simulation to show that our proposed approach can learn a good policy approximation by optimizing the global system EE in LoRa networks. Interesting directions of future work include integrating trajectory optimization for multiple flying LoRa gateways for efficient resource management. Our approach could be extended by exploring the Age of Information (AoI) scheme to enhance the freshness of data collection.